Solve the system of equation
-x+3y=9
3x-2y=4

Answer by checkley77(12844)

-x+3y=9 multipl this equation by 3 & add.
3x-2y=4
-3x+9y=27
--------------------
7y=31
y=31/7 ans.
-x+3*31/7=9
-x+93/7=9
-x=9-93/7
-x=(9*7-93)/7
-x=(63-93)/7
-x=-30/7
x=30/7 ans.
Proof:
3*30/7-2*31/7=4
90/7-62/7=4
28/7=4
4=4